# A Typology of Synthetic Datasets for Dialogue Processing in Clinical Contexts

### v1

@@ Abstract
Synthetic data sets are used across linguistic domains and NLP tasks, particularly in scenarios where authentic data is limited (or even non-existent). One such domain is that of clinical (healthcare) contexts, where there exist significant and long-standing challenges (e.g., privacy, anonymization, and data governance) which have led to the development of an increasing number of synthetic datasets. One increasingly important category of clinical dataset is that of clinical dialogues which are especially sensitive and difficult to collect, and as such are commonly synthesized.
While such synthetic datasets have been shown to be sufficient in some situations, little theory exists to inform how they may be best used and generalized to new applications. In this paper, we provide an overview of how synthetic datasets are created, evaluated and being used for dialogue related tasks in the medical domain. Additionally, we propose a novel typology for use in classifying types and degrees of data synthesis, to facilitate comparison and evaluation.
@@ A Typology of Synthetic Datasets for
Steven Bedrick Division of Informatics & Clinical Epidemiology Oregon Health and Science University bedricks@ohsu.edu                        A. Seza Doğruöz LT3, IDLab Universiteit Gent as.dogruoz@ugent.be                        Sergiu Nisioi HLT Research Center University of Bucharest sergiu.nisioi@unibuc.ro
1 Introduction
Synthetic datasets are increasingly used across linguistic domains and NLP tasks, particularly in scenarios where authentic data is limited or even non-existent. One such domain is that of clinical language, where there exist significant and long-standing challenges relating to privacy, anonymization, and data governance, which have led to the development of an increasing number of synthetic datasets Quintana (2020); Long et al. (2024).
A particularly fast-growing and relevant domain of clinical language is that of the “clinical dialogue,” i.e., a dialogue between a patient and a provider (sometimes referred to as a clinical or medical conversation). NLP tools to process such language are seeing rapid adoption in practice, most notably in the form of tools for automated generation of clinical documentation (Tierney et al., 2024; Finley et al., 2018) as well as for more traditional dialogue applications (e.g., clinical chatbots). Such dialogues may consist of spoken language, or may be written (taking place via an online discussion forum, a chat-style interface, etc.).
@@ 1 Introduction
While synthetic datasets have been shown to be sufficient in some situations, little theory exists to inform how they may be best used and generalized to new NLP applications. Towards that end, our goals in this paper are 1) to provide an overview of how synthetic data sets are created, evaluated and used for dialogue-related tasks in the medical domain via a literature review, and 2) classify these data sets according to a novel typology.
From an initial total of 1,626 papers (679 from PubMed, 591 from ACL Anthology, and 356 from dblp), successive filtering steps—including LLM-assisted classification and manual review—reduced the pool to 25 included papers (4 from PubMed, 11 from ACL Anthology, and 10 from dblp). After removing overlapping entries, the total number of unique papers is 20 (see Table 2 in Appendix C).
@@
One of the challenges toward this goal is a lack of consensus on what is meant by the word “synthetic” in terms of datasets on dialogues. The canonical definition of synthetic data is that of the U.S. Census Bureau (2025), which may be summarized as follows: given a dataset $D$, generate a new dataset $D^{\prime}$ in a way that a) individual instances in $D^{\prime}$ are not present in $D$, b) the relevant statistical properties of $D^{\prime}$ match those of $D$, and c) original instances from $D$ may not be recovered from $D^{\prime}$. [^1] This formulation echos that proposed in the earliest literature on synthetic microdata (Rubin, 1993). [^2] Throughout this paper, we use “microdata” in the statistical sense of the term; i.e., referring to the individual data points contained in the dataset: sentences in a corpus, images in a collection, responses from a survey, etc. This definition fits well in the context of a dataset consisting of structured data. However, as noted out by Aggarwal and Yu (2008), it is more challenging to apply this to textual data, for three primary reasons.
@@
First, as compared with traditional structured data, the processes to generate novel linguistic data points are far more complex. It is relatively straightforward to generate synthetic values for scalar or categorical variables by identifying appropriate distributions and then sampling using standard numerical techniques. However, producing language that preserves the semantic and discursive properties of the original necessitates much more sophisticated modeling and sampling techniques. Secondly (and relatedly), in the context of a structured data set, the phrase “relevant statistical properties” can be operationalized in terms of univariate or joint probabilities of the constituent variables. However, for linguistic data, the possibility space is far larger, and relying on simple statistical methods (e.g. perplexity) is unlikely to have satisfactory results. That is to say, the final dataset is unlikely to be truly “equivalent” to the original in linguistic terms (see §3). Finally, the question of re-identifiability (constraint “c” from the above definition) is more subtle and complex to assess in the context of linguistic data.
@@
Another challenging aspect of synthetic data is that artificiality is not a binary property: a dataset may be “more” or “less” synthetic, depending on its construction and the point of view of evaluation. All datasets, being curated and abstracted representations of some sampled reality, are inherently synthetic; [^3] See Appendix A for a more thoroughly developed definition of what constitutes a “dataset.” this process necessarily involves intervention around questions of inclusion/exclusion, representation, and structure. As Bowker (2005) put it, “raw data is both an oxymoron and a bad idea; to the contrary, data should be cooked with care.” However, while all datasets may be synthetic, some datasets are more synthetic than others. Literature on synthetic data distinguishes between datasets that are “fully synthetic,” being those that contain no original, genuine microdata, and “partially synthetic,” in which only certain variables are perturbed Reiter (2003). In terms of structured medical data, Gonzales et al. (2023) and Murtaza et al. (2023) provide comprehensive discussions on this and other relevant taxonomies. How might this framing, designed for structured datasetes consisting of discrete numerical variables, apply to linguistic data, in which we are less concerned with statistical properties of discrete variables and more with the semantic and syntactic properties of language?
@@
Exploring this question is the overall goal of this position paper. First, we review various ways in which synthetic language datasets are produced (§2). Next, we more formally define our terms to explore this continuum of synthesis (§3), and propose a typology that captures these subtleties in a practical and actionably useful way (§3.1). We also report the results of a literature review of datasets relating to clinical dialogues (§4.1) and use our proposed typology to describe existing datasets. However, we aim for a position paper (not a survey). Finally, we close with a discussion that extends beyond the microdata-centric view of our paper, to explore how realism and synthesis may relate to contextual aspects of a dataset (§4.2).
2 Means of Production
@@
Synthetic data production methods are frequently categorized into “process-driven” [^4] Sometimes referred to as “theory-” or “knowledge-driven.” and “data-driven” methods. Under the former, whatever generative process is carrying out the synthesis is informed by human knowledge and/or by information that is exogenous to the dataset itself. The latter describes methods that use properties of the dataset itself to derive the generative model (Goncalves et al., 2020; Murtaza et al., 2023).
This framing emphasizes description of the mechanical process by which the synthesis is done. While important, it does not describe what is being synthesized, nor does it address the degree to which a given dataset should be considered synthetic.
@@
From our review, synthetic clinical dialogue datasets are created using one of several methods. First, they can be created by taking “genuine” data (i.e., records of authentic patient-provider conversations that actually occurred in the “real world”), and perturbing it in some way. This perturbation may be intended to obscure identifiable information about actual patients, to produce a larger volume of comparable data than originally existed (e.g. Modersohn et al. (2022)), or as a form of data augmentation (Abdollahi et al., 2021). Perturbation, due to its relative technical simplicity, has been widely used, particularly for purposes of de-identification (Negash et al., 2023).
@@
Second, they can be created de novo via a human-driven manual process. For example, in medical education, it is very common for human actors to collaborate with clinical trainers or trainees to simulate various kinds of clinical interactions Rutherford-Hemming et al. (2024). While it is not generally the goal of such activities when done in a didactic context, the resulting recorded simulated encounters, drafted notes, etc., can serve as rich sources of data that are free from concerns about patient privacy. Of course, this sort of simulation activity can also be carried out specifically for the purpose of generating a re-usable dataset (Fareez et al., 2022; Yim et al., 2023).
@@
Alternatively, human annotators (trained or otherwise) may be tasked with writing the entire end-to-end conversation given pre-existing parameters and guidelines (Shi et al., 2023; Ben Abacha et al., 2023). In such scenarios, the resulting dialogue is not the natural result of humans interacting. It is essentially “screenwritten,” with the quality and utility of the resulting artifact depending heavily on the knowledge, experience, and skill of the author.
@@
In simulated scenarios involving humans, the pragmatics and paralinguistic components of resulting dialogues may be less representative of what would be seen in a true naturalistic exchange. Furthermore, we would expect a “screenwritten” scenario imagined by a single author to be less representative than a simulation involving multiple humans interacting in a more naturalistic way. However, depending on the intended use of the resulting data, this may not be an issue if the primary concern is the semantic or lexical contents of the exchange, as opposed to the pragmatic or discursive contents. [^5] See section 4.2 for additional discussion on this point.
@@
Finally, simulated datasets may be created mechanically de novo using a natural language generation (NLG) model (Frei and Kramer, 2023). This may take the form of prompted generation (i.e., an LLM being prompted to write a note or simulate a dialogue given certain criteria) or in some cases by literally having two LLMs “talk” to one another in a simulated dialogue, in an effort to increase the variety of the resulting output (Gray et al., 2024).
@@
Human-simulated scenarios have a long history of use in medical education and evaluation (Issenberg et al., 1999; McGaghie et al., 2010), resulting in a large and well-validated family of methodologies designed to ensure that the resulting interactions are sufficiently representative for their intended purpose (typically clinical training and assessment). Despite early and longstanding interest (Starkweather et al., 1967), entirely de novo dialogue generation has been relatively rarely used in clinical education. However, modern systems are beginning to be evaluated for use in medical training (Awada et al., 2024; Holderried et al., 2024), leading to the inevitable creation of a hybrid form of synthetic dataset, in which a human role-plays either the patient or the provider against the LLM playing the opposite role Campillos-Llanos et al. (2021); Abercrombie and Rieser (2022); Zhang et al. (2023)
@@
Given this variety of means of generation, we argue that it is inaccurate to think of a dataset as being either “synthetic” or “real”. Instead, any given dataset exists as a point along a continuum between the two poles. After all, even the most “real-world” datasets are inevitably the result of numerous manual interventions and decisions (Gitelman, 2013; Passi and Sengers, 2020; Sambasivan et al., 2021).
@@
Similarly, an “anonymized” dataset is in some sense synthetic (the degree to which this is the case depends on the method of anonymization in question). At one end of the spectrum, imagine a scheme in which identifiers are replaced with a class token of some kind (e.g. PATIENT_NAME, DAY_OF_WEEK, etc.) in a purely deterministic and uniform manner. At the other end of that spectrum, consider a method in which identifiers are replaced with statistically-plausible alternative values (“pseudonymization”) [^6] See Uzuner et al. (2007) for a detailed description of such a method, as well as an overview on deidentification in non-dialogue clinical text in general.
@@
Clearly, in both cases, the process of applying such a method results in the creation of a new and distinct dataset. We argue, however, that the second case is “more synthetic” in that there are two levels of synthesis in play: first, the simple alteration of the original un-modified data, and second, in that some designed process informed the pseudonymization. This process must itself necessarily bring with it working assumptions, and may also depend on yet additional sources of data, further distancing the resulting synthetic dataset from whatever grounding in reality it may have had earlier in its development.
@@ 3 Let’s Get Synthetic
In response to the complexities surrounding the question of whether a given dataset is “synthetic” or not, we propose a conceptual framework whose aim is to provide structure and clarity when describing datasets that have a synthetic component, and to facilitate meaningful comparisons. For our purposes, we define a dataset as a collection of information-bearing objects that have been purposefully assembled or otherwise grouped, with the goal of serving a particular representational function, with the intention of being used in discourse or analysis of some kind. See Appendix A for a detailed discussion of this definition and its constituent elements.
@@
How does our above definition fit with synthetic data? A dataset is per definition a human-generated artifact. At all stages of collection, processing, and inclusion, choices are made that separate the resulting dataset from its originating context. However, we first define a “naturalistic” dataset as one that fulfills the terms of our above definition, along with the additional constraint that the microdata contained in the dataset are a) intended by the designers to be faithful representations of some genuine originating phenomenon, and b) are understood to be such by the users of the dataset.
@@
For example, consider the MedDialog corpus of patient-provider dialogues taken from an online medical consultation website (Zeng et al., 2020). The designers of this dataset understood and believed these dialogues to represent exchanges that actually took place between human patients and human providers who actually existed. While they may have applied criteria to which dialogues were included (perhaps focusing on a particular category of patient, or clinical specialty), the lexical, semantic, and pragmatic properties of the language were left un-modified in the hope of comprising a maximally representative sample. Consumers of this dataset would similarly choose to use the dataset with the understanding that its contents were representative of exchanges that actually took place.
@@
In comparison, consider the aci subset of the Aci-bench dataset (Yim et al., 2023), which consists of transcribed exchanges between an actual doctor and a volunteer “patient,” acting out a scenario based on symptom-driven prompts. The exchanges in such a dataset are “real” in the sense that they did, in fact, take place, and occurred between actual humans. However, they are not “real” in that they arose from an artificial scenario, as opposed from the actual real-world domain that they are attempting to represent.
@@
The creators of Aci-bench, of course, attempted to design their scenarios to be as “realistic” as possible, and gave careful consideration to their design. Their goal was to create a dataset whose synthetic features did not affect those aspects of its form or content to an extent that compromised the dataset’s fundamental validity [^7] Here we use “validity” in the more specific senses of construct and measurement validity; see Brewer and Crano (2014) for an overview of these terms, and Jacobs and Wallach (2021) for a detailed and enlightening discussion of how these concepts apply in machine learning research. for the analyses of the sort that the dataset was intended by its authors and users to support. Decisions about whether, when, and how to use such a dataset typically hinge on the extent to which this assumption holds for one’s particular purpose.
@@
Finally, consider the DoPaCo dataset described by Chen et al. (2023a). Here, the authors began with a dataset of transcribed and de-identified genuine doctor-patient conversations, to which they had extremely restricted and time-limited access, and which could not be shared or re-distributed. They used this corpus to fine-tune a pre-trained dialogue-generation Transformer model, and then used this model to generate a large corpus of synthetic dialogues, which they then used for the rest of their work. One may argue that such a dataset is in some sense meaningfully or usefully representative of a “real” dataset of doctor-patient conversations, according to any number of rubrics. [^8] Given that it was produced via a sequence-to-sequence model, one may accurately and literally describe it as consisting of a set of samples drawn from a statistical model of the originating “real” dataset; as such, there exist numerous statistical tools that we may use to quantify the degree to which such samples are or are not representative, in a purely statistical sense, of the originating data. Perplexity and KL-Divergence are two such methods of quantifying this statistical relationship. However, we note that there is a long-standing and robust literature (Galliers and Spärck Jones, 1993; Iyer et al., 1997; Ito et al., 1999) on the limitations of such purely statistical measures in terms of their external and ecological validity (Brewer and Crano, 2014; Egami and Hartman, 2023). But, it is also equally true that such a dataset, however statistically similar it may be to its original training data, is not comprised of “real” dialogue data in the same sense as Aci-bench or MedDialog.
@@
This question of “real-ness” is an entirely separate question from whether such a dataset is useful or sufficient for a given purpose. We are again using “real” as a descriptive rather than a normative term, and of course even highly synthetic datasets have many uses. As such, we are emphatically not arguing that datasets that are more naturalistic are inherently superior in some way to datasets that are more synthetic. There may indeed exist scenarios where this is the case. However, that question is orthogonal to our purpose in the present work. Finally, we also note that many of the synthetic datasets we describe in this paper are accompanied by more or less detailed evaluations by their creators, seeking to investigate the degree to which they are or are not sufficient for specific applications.
3.1 A Typology for Human/Machine Microdata Interventions
We propose a three-level typology of “types” of intentional intervention to the content of a dataset’s microdata (illustrated in Table 1). Our typology’s categories are inspired by the perturbation-generation dichotomy described by Domingo-Ferrer (2008), adapted and expanded in terms appropriate to the complexities of linguistic datasets (as opposed to traditional statistical datasets). Given the wide diversity in ways datasets are synthesized, our categorization schema provides a practical balance between expressivity and simplicity.

### v2

@@ Abstract
Synthetic data sets are used across linguistic domains and NLP tasks, particularly in scenarios where authentic data is limited (or even non-existent). One such domain is that of clinical (healthcare) contexts, where there exist significant and long-standing challenges (e.g., privacy, anonymization, and data governance) which have led to the development of an increasing number of synthetic datasets. One increasingly important category of clinical dataset is that of clinical dialogues which are especially sensitive and difficult to collect. Therefore, they are commonly synthesized. While such synthetic datasets have been shown to be sufficient in some situations, little theory exists to inform how they may be best used and generalized to new applications. In this paper, we provide an overview of how synthetic datasets are created, evaluated and used for dialogue related tasks in the medical domain. Additionally, we propose a novel typology for use in classifying types and degrees of data synthesis, to facilitate comparison and evaluation.
Keywords: health care, synthetic data, clinical conversation datasets
[…]
@@ A Typology of Synthetic Datasets for
Steven Bedrick   A. Seza Doğruöz   Sergiu Nisioi
Division of Informatics & Clinical Epidemiology
 Oregon Health & Science University
LT3, IDLab, Universiteit Gent
HLT Research Center, University of Bucharest
bedricks@ohsu.edu, as.dogruoz@ugent.be, sergiu.nisioi@unibuc.ro
Abstract content
1.  Introduction
Synthetic datasets are increasingly used across linguistic domains and NLP tasks, particularly in scenarios where authentic data is limited (or even non-existent). One such domain is that of clinical language, where there are significant and long-standing challenges relating to privacy, anonymization, and data governance, which have led to the development of an increasing number of synthetic datasets quintana2020; long-etal-2024-llms.
A particularly fast-growing and relevant domain of clinical language is that of the “clinical dialogue (i.e., a dialogue between a patient and a health provider which can also be referred to as a clinical or medical conversation). NLP tools to process such language are seeing rapid adoption in practice, most notably in the form of tools for automated generation of clinical documentation (Tierney:2024; Finley:2018) and for more traditional dialogue applications (e.g., clinical chatbots). Such dialogues may occur via spoken language, or may be written (taking place via an online discussion forum, a chat-style interface, etc.).
@@ 1.  Introduction
While synthetic datasets have been shown to be sufficient in some situations, little theory exists to inform the scientific audience about how these datasets may be best used and generalized to new NLP applications. Towards that end, our goals in this paper are 1) to provide an overview of how synthetic data sets are created, evaluated and used for dialogue-related tasks in the medical domain via a literature review, and 2) to classify these data sets according to a novel typology.
@@
One of the challenges toward this goal is a lack of consensus on what is meant by the word “synthetic” in terms of datasets on dialogues. The canonical definition of synthetic data is that of the census:2025aa, which may be summarized as follows: given a dataset $D$, generate a new dataset $D^{\prime}$ in a way that a) individual instances in $D^{\prime}$ are not present in $D$, b) the relevant statistical properties of $D^{\prime}$ match those of $D$, and c) original instances from $D$ may not be recovered from $D^{\prime}$. [^1] This formulation echos that proposed in the earliest literature on synthetic microdata (DonaldB.Rubin:1993). [^2] Throughout this paper, we use “microdata” in the statistical sense of the term; i.e., referring to the individual data points contained in the dataset: sentences in a corpus, images in a collection, responses from a survey, etc. This definition fits well in the context of a dataset consisting of structured data. However, as noted out by aggarwalGeneralSurveyPrivacyPreserving2008, it is more challenging to apply this to textual data, for three primary reasons.
@@
First, as compared with traditional structured data, the processes to generate novel linguistic data points are far more complex. It is relatively straightforward to generate synthetic values for scalar or categorical variables by identifying appropriate distributions and then sampling using standard numerical techniques. However, producing language that preserves the semantic and discursive properties of the original necessitates more sophisticated modeling and sampling techniques. Secondly, in the context of a structured data set, the phrase “relevant statistical properties” can be operationalized in terms of univariate or joint probabilities of the constituent variables. However, for linguistic data, the possibility space is far larger, and relying on simple statistical methods (e.g. perplexity) is unlikely to have satisfactory results. That is to say, the final dataset is unlikely to be truly “equivalent” to the original in linguistic terms (see §3). Finally, the question of re-identifiability (constraint “c” from the above definition) is subtle and complex to assess in the context of linguistic data.
@@
Another challenging aspect of synthetic data is that artificiality is not a binary property. In that sense, a dataset may be “more” or “less” synthetic, depending on its construction and the point of view of evaluation. All datasets, being curated and abstracted representations of some sampled reality, are inherently synthetic to some degree; this process necessarily involves intervention around questions of inclusion/exclusion, representation, and structure. As bowkerMemoryPracticesSciences2005 put it, “raw data is both an oxymoron and a bad idea; to the contrary, data should be cooked with care.” However, while all datasets may be synthetic, some datasets are more synthetic than others. Literature on synthetic data distinguishes between datasets that are “fully synthetic,” being those that contain no original, genuine microdata, and “partially synthetic,” in which only certain variables are perturbed reiterInferencePartiallySynthetic2003. In terms of structured medical data, Gonzales:2023 and Murtaza:2023 provide comprehensive discussions on this and other relevant taxonomies. How can this framework, originally designed for numerical data, be applied to linguistic data, which focuses on meaning and grammar rather than statistics?
@@
Exploring this question is the overall goal of our paper. In line with this goal, first, we review various ways in which synthetic language datasets are produced (§2). Next, we more formally define our terms to explore this continuum of synthesis (§3), and propose a typology that captures these subtleties in a practical and actionably useful way (§3.1). We also report the results of a literature review of datasets relating to clinical dialogues (§4.1) and demonstrate how our proposed typology describes existing datasets. Finally, we close with a discussion that extends beyond the microdata-centric view of our paper, to explore how realism and synthesis may relate to contextual aspects of a dataset (§4.2).
2.  Means of Production
@@
Synthetic data production methods are frequently categorized into “process-driven” [^3] Sometimes referred to as “theory-” or “knowledge-driven.” and “data-driven” methods. Under the former, whatever generative process is carrying out the synthesis is informed by human knowledge and/or by information that is exogenous to the dataset itself. The latter describes methods that use properties of the dataset itself to derive the generative model (Goncalves:2020a; Murtaza:2023).
This framing emphasizes description of the mechanical process by which the synthesis is done. While important, it does not describe what is being synthesized, nor does it address the degree to which a given dataset should be considered synthetic. This omission is problematic when considering synthetic linguistic data, as there are numerous aspects of the data that might (or might not) be synthesized: topic, genre/register, grammatical or discursive structure, specific aspects of semantic content, individual lexical choices, etc. A given dataset might involve synthesis along one of these domains, while leaving other domains un-perturbed. To illustrate this issue, consider the question of how synthetic clinical datasets are created.
@@
From our review, we find that such datasets are created using one of several methods. First, they can be created by taking “genuine” data (i.e., records of authentic patient-provider conversations that actually occurred in the “real world”), and perturbing it in some way. This perturbation may be intended to obscure identifiable information about actual patients (Negash:2023), to produce a larger volume of comparable data than originally existed (e.g. modersohnGRASCCOFirstPublicly2022), or as a form of data augmentation (Abdollahi:2021).
@@
Second, a conversational dataset can be created de novo via a human-driven manual process. In medical education, it is very common for human actors to collaborate with clinical trainers or trainees to simulate various kinds of clinical interactions Rutherford-Hemming:2024. Recordings of such simulated encounters, as well as other artifacts such as encounter notes, etc., can serve as rich sources of data that are free from concerns about patient privacy. While such secondary use is rarely the goal of simulations when done in didactic contexts, they can also be performed specifically for the purpose of generating a re-usable dataset (fareez2022dataset; yim2023aci).
@@
Alternatively, human annotators (trained or otherwise) may be tasked with writing an entire end-to-end conversation given pre-existing parameters and guidelines (shi-etal-2023-midmed; ben-abacha-etal-2023-empirical). In such scenarios, the resulting dialogue is not the natural result of humans interacting. It is essentially “screenwritten,” with the quality and utility of the resulting artifact depending heavily on the knowledge, experience, and skill of the author.
@@
In simulated scenarios involving humans, the pragmatics and paralinguistic components of resulting dialogues may be less representative of what would be seen in a true naturalistic exchange. Similarly, we would expect a “screenwritten” scenario imagined by a single author to be less representative than a simulation involving multiple humans interacting in a more naturalistic way. However, depending on the intended use of the resulting data, this may not be an issue: for example, if the primary content of interest is the semantic or lexical contents of the exchange, as opposed to the pragmatic or discursive contents. [^4] See section 4.2 for additional discussion.
@@
Finally, simulated datasets may be created mechanically de novo using a natural language generation (NLG) model (frei2023annotated). This may take the form of prompted generation (i.e., an LLM being prompted to write a note or simulate a dialogue given certain criteria) or in some cases by literally having two LLMs “talk” to one another in a simulated dialogue, in an effort to increase the variety of the resulting output (grayIncreasingRealismVariety2024).
@@
Human-simulated scenarios have a long history of use in medical education and evaluation (Issenberg:1999; McGaghie:2010), resulting in a large and well-validated family of methodologies designed to ensure that the resulting interactions are sufficiently representative for their intended purpose (typically clinical training and assessment). Despite early and longstanding interest (Starkweather:1967), computer-driven de novo dialogue generation has been relatively rarely used in clinical education. However, modern systems are beginning to be evaluated for use in medical training (Awada2024; Holderried:2024), leading to the inevitable creation of an additional form of synthetic dataset, in which a human role-plays either the patient or the provider against the LLM playing the opposite role campillos2021lessons; abercrombie-rieser-2022-risk; zhang-etal-2023-huatuogpt
@@
Given this variety of means of generation, we argue that it is inaccurate to think of a dataset as being either “synthetic” or “real”. Instead, we argue that any given dataset exists as a point along a continuum between the two poles. After all, even the most “real-world” datasets are inevitably the result of numerous manual interventions and decisions (gitelmanRawData2013; passiMakingDataScience2020; sambasivanEveryoneWantsModel2021).
@@
Similarly, we argue that an otherwise-genuine dataset that has been “anonymized” is, in some sense synthetic (the degree to which this is the case depends on the method of anonymization in question). At one end of the spectrum, imagine a scheme in which identifiers are replaced with a class token of some kind (e.g. PATIENT_NAME, DAY_OF_WEEK, etc.) in a purely deterministic and uniform manner. At the other end of that spectrum, consider a method in which identifiers are replaced with statistically-plausible alternative values (“pseudonymization”) [^5] See uzunerEvaluatingStateoftheartAutomatic2007 for a detailed description of such a method, as well as an overview on deidentification in non-dialogue clinical text in general.
@@
In both cases, the process of applying such a method results in the creation of a new and distinct dataset. We argue, however, that the second case is “more synthetic” in that there are two levels of synthesis in play: first, the simple alteration of the original un-modified data, and second, in that some designed process informed the pseudonymization. This process must itself necessarily bring with it working assumptions, and may also depend on yet additional sources of data, further distancing the resulting synthetic dataset from whatever grounding in reality it may have had earlier in its development.
@@ 3.  What Do We Mean By “Synthetic”?
In response to the complexities surrounding the question of whether a given dataset is “synthetic” or not, we propose a conceptual framework whose aim is to provide structure and clarity when describing datasets that have a synthetic component, and to facilitate meaningful comparisons. For our purposes, we define a dataset as a collection of information-bearing objects that have been purposefully assembled or otherwise grouped, with the goal of serving a particular representational function, with the intention of being used in analysis (broadly defined) of some kind.
@@
How does our above definition fit with synthetic data? A dataset is, per definition, a human-generated artifact. At all stages of collection, curation, and processing, choices are made that separate the resulting dataset from its originating context. As such, we further define a “naturalistic” dataset as one that fulfills the terms of our above definition, along with the additional constraint that the microdata contained in the dataset are a) intended by the designers to be faithful representations of some genuine originating phenomenon, and b) are understood to be such by the users of the dataset.
@@
For example, consider the MedDialog corpus of patient-provider dialogues taken from an online medical consultation website (zeng-etal-2020-meddialog). The designers of this dataset understood and believed these dialogues to represent exchanges that actually took place between human patients and human providers who actually existed. While they may have applied criteria to which dialogues were included (perhaps focusing on a particular category of patient, or clinical specialty), the lexical, semantic, and pragmatic properties of the language were left un-modified in the hope of comprising a maximally representative sample. In turn, downstream users of this dataset would understand its contents to be representative of exchanges that actually took place (i.e., to be “real messages”).
@@
In comparison, consider the aci subset of the Aci-bench dataset (yim2023aci), which consists of transcribed exchanges between an actual doctor and a volunteer “patient,” acting out a scenario based on symptom-driven prompts. The exchanges in such a dataset are “real” in the sense that they did, in fact, take place, and occurred between actual humans. However, they are not “real” in that they arose from an artificial scenario, as opposed from the actual real-world domain that they are attempting to represent.
@@
The creators of Aci-bench, of course, attempted to design their scenarios to be as “realistic” as possible, and gave careful consideration to their design. Their goal was to create a dataset whose synthetic features did not affect those aspects of its form or content to an extent that compromised the dataset’s fundamental validity [^6] Here we use “validity” in the specific senses of construct and measurement validity; see brewer_crano_2014 for an overview of these terms, and jacobs_wallach_2021 for a discussion of how these concepts apply in machine learning research. for the analyses of the sort that the dataset was intended by its authors and users to support. Decisions about whether, when, and how to use such a dataset typically hinge on the extent to which this assumption holds for one’s particular purpose.
@@
Finally, consider the DoPaCo dataset  (chen2023investigating). Here, the authors began with transcribed and de-identified genuine doctor-patient conversations, to which they had extremely restricted and time-limited access, and which could not be re-distributed. They used this corpus to fine-tune a pre-trained dialogue-generation Transformer model, and then used this model to generate a large corpus of synthetic dialogues, which they used for the rest of their work. One may argue that such a dataset is in some sense meaningfully or usefully representative of a “real” dataset of doctor-patient conversations, according to any number of rubrics. [^7] Given that it was produced via a sequence-to-sequence model, one may accurately and literally describe it as consisting of a set of samples drawn from a statistical model of the originating “real” dataset; as such, there exist statistical tools that we may use to quantify the degree to which such samples are or are not representative (in a purely statistical sense) of the originating data, such as Perplexity and KL-Divergence. However, we note that there is a long-standing and robust literature (Galliers:1993; iyer_1997; ito_1999) on the limitations of such purely statistical measures in terms of their external and ecological validity (brewer_crano_2014; Egami:2023). But, it is also equally true that such a dataset, however statistically similar it may be to its original training data, is not comprised of “real” dialogue data in the same sense as Aci-bench or MedDialog. Rather than representing true discursive choices made by humans interacting with one another, the dialogues in the DoPaCo dataset reflect that model’s statistical approximation of such choices, as represented its training data.
@@
This question of “real-ness” is an entirely separate question from whether such a dataset is useful or sufficient for a given purpose. We are again using “real” as a descriptive rather than a normative term, and of course even highly synthetic datasets have many uses. As such, we are emphatically not arguing that datasets that are more naturalistic are inherently superior in some way to datasets that are more synthetic, though there may indeed exist scenarios where this is the case. However, that question is orthogonal to our purpose in the present work. Finally, we note that the synthetic datasets we describe in this paper are accompanied by more or less detailed evaluations by their creators, seeking to investigate the degree to which they are or are not sufficient for specific applications.
3.1.  A Typology for Human/Machine Microdata Interventions
We propose a three-level typology of “types” of intentional intervention to the content of a dataset’s microdata (illustrated in Table 1). Our typology’s categories are inspired by the perturbation-generation dichotomy described by domingo-ferrerSurveyInferenceControl2008, adapted and expanded in terms appropriate to the complexities of linguistic datasets (as opposed to traditional statistical datasets). Additionally, we observe that the interventions involved in producing a synthetic linguistic dataset may be carried out by either humans or machines (or both) during the course of a dataset’s creation. In practice, many synthetic datasets involve a mix of both, and adequately capturing the ways in which a given dataset is “synthetic” requires that both aspects be described. As such, our typology involves categorizing a dataset along dimensions of both the human and machine involvement (Table 1). This increased expressiveness further helps breakdown the false binary of “synthetic” vs. “real,” and allows clearer descriptions of a given dataset.